A speech class has seven freshmen and five sophomores. Three people are selected at random to give speeches in sequence. Find the probability of each of the following. (Round your answers to four decimal places.)
(a) the first is a sophomore, the second is a freshman, and the third is a sophomore
(b) all three are freshmen
(c) the first is a sophomore and the other two are freshmen

Answer by mathmate(429)

Question:
A speech class has seven freshmen and five sophomores. Three people are selected at random to give speeches in sequence. Find the probability of each of the following. (Round your answers to four decimal places.)
(a) the first is a sophomore, the second is a freshman, and the third is a sophomore
(b) all three are freshmen
(c) the first is a sophomore and the other two are freshmen

Solution:
Use multiplication rule
7 freshmen+5 sophomores = 12 students.
(a) SFS
P(SFS)=5/12*(7/11)*(4/10)=7/66
(b) FFF
P(FFF)=7/12*6/11*5/10=7/44
(c) SFF
P(SFF)=5/12*7/11*6/10=7/44